Melissa's cat is five years younger than she is. In two years,Melissa will be twice as old as her cat.How old are they now?
Answer by checkley77(12844)

Melissa age now=x
Cat's age now (x-5)
In 2 years Melissa will be (x+2)
The cat will be (x-5+2) or (x-3)
(x+2)=2(x-3)
x+2=2x-6
x-2x=-6-2
-x=-8
x=8 Melissa's age now
8-5=3 The cat's age now.
Proof:
8+2=2(8-3)
10=2*5
10=10